# PPR: Enhancing Dodging Attacks while Maintaining Impersonation Attacks on Face Recognition Systems

### v2

@@ 3.3 Pre-training Pruning Restoration Attack
where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, hight, width of the face images, respectively.
Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measures the ratio of adversarial perturbations to be set into zero. Let $s$ be the number of adversarial perturbation elements:
$s={\rm CHW}$ (9)
We arrange the elements in flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
@@
Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned. Given the importance calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
@@
where the colon denotes the slice operation to obtain the first $\kappa s$ elements.
The mask for pruning can be obtained using $\mathcal{W}$:
@@
Once the adversarial face examples are pruned, we can proceed to restore the adversarial perturbations in the pruned regions. The mask representing the regions for restoring the adversarial examples can be denoted as:
@@
For both the pre-training and restoration stages, various adversarial attacks can be employed. In the following, we will illustrate the pre-training and restoration stages of of our proposed method by using the Lagrangian attack as an example. In the stage of pre-training, we utilize Equation 7 as the pre-training loss $\mathcal{L}^{p}$ to craft the adversarial face examples:
$x^{adv}_{t}=x^{adv}_{t-1}-\beta{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{p}\right)$ (15)
where $t$ is the iteration of optimization process of adversarial examples, $\beta$ is the step size when optimizing the adversarial face examples in the pre-training stage.
Once the adversarial examples are pre-trained, we utilize our proposed method of pruning adversarial face examples to free up some adversarial perturbations. During the restoration stage, we restore the adversarial face examples in the previously pruned region using the following loss function:
$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (16)
where $\tilde{\lambda}$ is a weight which is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks.
@@
$x^{adv}_{t}=x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)$ (17)
@@
where $\gamma$ is the step size when optimizing the adversarial face examples in the restoration stage. The algorithm of our proposed method based on Lagrangian attack are illustrated in Algorithm 1.
Algorithm 1 Pre-training Pruning Restoration Attack Based on Lagrangian
@@
0:  Negative face image pair $\{x^{s},x^{t}\}$, the step size of the adversarial perturbations $\beta$, the number of the iterations of pre-training stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$.
@@
3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Pre-training Stage
@@
4 Experiments
4.1 Experimental Setting
Datasets. We have opted to use the LFW dataset (Huang et al., 2007) and the CelebA-HQ dataset (Karras et al., 2018) for our experiments. LFW serves as an unconstrained face dataset for FR, while CelebA-HQ consists of high-quality images. These two datasets are widely employed in research on adversarial attacks in FR (Xiao et al., 2021; Yin et al., 2021; Hu et al., 2022). The LFW and CelebA-HQ datasets utilized in our experiments are identical to those employed in previous studies (Zhou et al., 2023b, a).
Table 1: Comparisons of dodging ASR results for attacks on the LFW and CelebA-HQ datasets. The surrogate models are presented in the first column, and the victim models are listed in the second row.
LFW CelebA-HQ
Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
IR152 DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
SSA 78.8 / 100.0 5.7 / 22.5 0.9 / 10.6 83.8 / 99.9 7.2 / 19.6 0.4 / 5.6
DFANet 98.9 / 100.0 1.4 / 4.2 0.0 / 0.3 98.9 / 100.0 2.3 / 6.0 0.0 / 0.4
SIA 81.7 / 100.0 13.0 / 37.5 0.8 / 8.9 78.4 / 100.0 13.2 / 35.6 0.7 / 7.4
BSR 52.4 / 100.0 5.3 / 17.6 0.1 / 1.5 48.5 / 99.9 5.4 / 18.0 0.3 / 1.9
BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
FaceNet DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
SSA 6.0 / 14.0 97.5 / 99.9 6.6 / 26.2 2.0 / 5.5 96.9 / 99.7 4.2 / 14.6
DFANet 1.6 / 3.3 99.8 / 99.9 0.4 / 2.7 0.5 / 2.6 99.1 / 100.0 0.8 / 4.1
SIA 11.2 / 20.6 99.5 / 99.9 8.7 / 21.2 4.0 / 8.9 99.4 / 99.9 5.4 / 13.7
BSR 12.2 / 19.2 98.6 / 99.9 9.0 / 17.8 4.6 / 10.1 98.8 / 99.9 5.3 / 14.1
BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
MF DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
SSA 0.7 / 4.1 6.1 / 23.5 98.3 / 100.0 0.0 / 0.6 3.9 / 18.5 93.3 / 100.0
DFANet 0.2 / 1.0 1.5 / 5.8 99.6 / 100.0 0.0 / 0.2 1.1 / 7.7 99.1 / 100.0
SIA 1.0 / 5.9 10.6 / 36.6 98.4 / 100.0 0.1 / 2.4 9.0 / 24.4 96.3 / 100.0
BSR 0.4 / 1.5 3.7 / 14.7 84.9 / 100.0 0.1 / 0.6 2.9 / 12.6 77.6 / 100.0
BPFA 0.9 / 4.1 4.6 / 20.4 97.7 / 100.0 0.0 / 2.3 4.0 / 20.4 96.2 / 100.0

### v3

@@ 3.3 Adversarial Pruning Attack
where $\mathcal{I}\in\mathbb{R}^{\rm CHW}$. ${\rm C}$, ${\rm H}$, and ${\rm W}$ are the channel number, height, and width of the face images, respectively.
Once the priority values of the adversarial perturbations are quantified, we employ these values to release less prioritized adversarial perturbations. Let $\kappa$ be the sparsity ratio for pruning the adversarial face examples that measure the ratio of perturbations to be set into zero. Let $s={\rm CHW}$ be the number of adversarial perturbation elements. We arrange the elements in a flattened vector of $\mathcal{I}$ in ascending order (from the lowest to the highest):
@@
Let $\mathcal{W}$ be the set of the elements of the adversarial perturbations to be pruned. Given the priority calculation method for pruning, the value of $\mathcal{W}$ can be calculated as follows:
@@
where the colon denotes the slice operation to obtain the first $\kappa s$ elements. The pruning mask, which has the same shape as $\mathcal{I}$, can be obtained by utilizing $\mathcal{W}$:
@@
where $\bar{x}^{adv}$ is the adversarial face example after pruning.
Biased Gradient Adaptation. During the Restoration stage, we restore the adversarial face examples in the previously pruned region using our proposed Biased Gradient Adaptation. Biased Gradient Adaptation using the following loss function to craft gradient biased to the dodging attacks to adapt the crafted adversarial examples into the space that favors dodging attacks.
$\mathcal{L}^{r}=\tilde{\lambda}\mathcal{L}^{i}+\mathcal{L}^{d}$ (10)
where $\tilde{\lambda}$ is a weight that is lower than $\lambda$ that is objective for crafting adversarial face examples that favor dodging attacks. The mask representing the regions for restoring the adversarial examples can be denoted as:
@@
Subsequently, we utilize the following formula to restore the pruned adversarial face examples:
@@
$\begin{split}x^{adv}_{t}=\prod\limits_{x^{s},\epsilon}\left(x^{adv}_{n}+\mathcal{A}\odot\left(x^{adv}_{t-1}-\gamma{\rm sign}\left(\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}\right)-\bar{x}^{adv}\right)\right)\end{split}$ (12)
@@
where $\gamma$ is the step size when optimizing the adversarial face examples in the Restoration stage, $\bar{x}^{adv}$ is the adversarial example crafted by the Priming Stage, and $\nabla_{x^{adv}_{t-1}}\mathcal{L}^{r}$ is the biased gradient. The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in the appendices.
4 Experiments
4.1 Experimental Setting
Datasets. We opt to use the LFW [20], CelebA-HQ [24], and FFHQ [25] datasets for our experiments. LFW serves as an unconstrained face dataset for FR. CelebA-HQ and FFHQ consists of high-quality images. The LFW and CelebA-HQ utilized in our experiments are identical to those employed in [72; 71], while FFHQ is the corresponding dataset provided by the Sibling-Attack official page, ensuring the consistency for analysis.
Figure 3: The ASR on FR models trained by multiple algorithms.
Figure 4: Comparisons of ASR (%) on LFW with adversarial robust models as victim models.
Table 1: Comparisons of dodging ASR (%) results for attacks on the LFW and CelebA-HQ datasets. The surrogate models are presented in the first column, and the victim models are listed in the second row. The numbers before and after the slash represent the results of the baseline attack and the attack that combines the baseline attack with our proposed attack, respectively.
LFW CelebA-HQ
Surrogate Model Attack IR152 FaceNet MF IR152 FaceNet MF
IR152 [17] DI 95.4 / 100.0 5.8 / 11.3 0.2 / 0.8 87.9 / 100.0 8.4 / 16.4 0.4 / 1.2
VMI 92.7 / 100.0 17.3 / 32.5 1.2 / 9.5 92.2 / 100.0 14.3 / 27.9 1.2 / 4.1
SSA 78.8 / 100.0 5.7 / 22.5 0.9 / 10.6 83.8 / 99.9 7.2 / 19.6 0.4 / 5.6
DFANet 98.9 / 100.0 1.4 / 4.2 0.0 / 0.3 98.9 / 100.0 2.3 / 6.0 0.0 / 0.4
SIA 81.7 / 100.0 13.0 / 37.5 0.8 / 8.9 78.4 / 100.0 13.2 / 35.6 0.7 / 7.4
BSR 52.4 / 100.0 5.3 / 17.6 0.1 / 1.5 48.5 / 99.9 5.4 / 18.0 0.3 / 1.9
BPFA 92.6 / 100.0 1.7 / 7.3 0.0 / 1.2 90.4 / 100.0 2.1 / 8.1 0.1 / 0.8
FaceNet [41] DI 5.3 / 10.3 99.8 / 99.9 3.1 / 10.3 1.5 / 3.1 99.4 / 99.9 1.8 / 4.7
VMI 9.7 / 14.3 99.8 / 99.9 6.2 / 13.2 3.1 / 7.1 99.3 / 99.8 3.6 / 9.3
SSA 6.0 / 14.0 97.5 / 99.9 6.6 / 26.2 2.0 / 5.5 96.9 / 99.7 4.2 / 14.6
DFANet 1.6 / 3.3 99.8 / 99.9 0.4 / 2.7 0.5 / 2.6 99.1 / 100.0 0.8 / 4.1
SIA 11.2 / 20.6 99.5 / 99.9 8.7 / 21.2 4.0 / 8.9 99.4 / 99.9 5.4 / 13.7
BSR 12.2 / 19.2 98.6 / 99.9 9.0 / 17.8 4.6 / 10.1 98.8 / 99.9 5.3 / 14.1
BPFA 4.7 / 16.8 98.6 / 100.0 1.6 / 15.0 1.1 / 4.2 99.0 / 100.0 0.6 / 5.1
MF [9] DI 2.2 / 7.3 18.2 / 36.4 99.2 / 100.0 0.1 / 2.5 12.1 / 31.3 95.2 / 100.0
VMI 1.0 / 2.8 8.4 / 20.9 99.7 / 100.0 0.2 / 0.4 5.2 / 15.0 98.2 / 100.0
SSA 0.7 / 4.1 6.1 / 23.5 98.3 / 100.0 0.0 / 0.6 3.9 / 18.5 93.3 / 100.0
DFANet 0.2 / 1.0 1.5 / 5.8 99.6 / 100.0 0.0 / 0.2 1.1 / 7.7 99.1 / 100.0
SIA 1.0 / 5.9 10.6 / 36.6 98.4 / 100.0 0.1 / 2.4 9.0 / 24.4 96.3 / 100.0
BSR 0.4 / 1.5 3.7 / 14.7 84.9 / 100.0 0.1 / 0.6 2.9 / 12.6 77.6 / 100.0
BPFA 0.9 / 4.1 4.6 / 20.4 97.7 / 100.0 0.0 / 2.3 4.0 / 20.4 96.2 / 100.0
Face Recognition Models. The normal trained FR models employed in our experiments include IR152 [17], FaceNet [41], MobileFace (abbreviated as MF) [9], ArcFace [9], CircleLoss [45], CurricularFace [22], MagFace [36], MV-Softmax [52], and NPCFace [67]. IR152, FaceNet, and MF are identical to those used in [64; 19; 72; 71]. ArcFace, CircleLoss, CurricularFace, MagFace, MV-Softmax, and NPCFace are the official models available in FaceX-ZOO [50]. Additionally, we incorporate adversarial robust FR models in our experiments, denoted as IR152adv, FaceNetadv, and MFadv, which are identical to those used in [72]. For calculating the ASR in impersonation and dodging attacks, we choose the thresholds based on FAR@0.001 on the entire LFW dataset.
Attack Setting. Without any particular emphasis, we set the maximum allowable perturbation magnitude to 10 based on the $L_{\infty}$ norm bound and utilize the Lagrangian attack method as the attack in both the Priming and Restoration stages. Additionally, we specify the maximum number of iterative steps as 200. For both the Priming and Restoration stages, the step size is uniformly designated as 1.0.
Evaluation Metrics. We employ Attack Success Rate (ASR) to evaluate the performance of various attacks. ASR signifies the proportion of successfully attacked adversarial examples out of all the adversarial examples. We use ASRi and ASRd to denote impersonation and dodging ASR, respectively. The detailed calculation methods for ASRi and ASRd are provided in the appendices.
Compared methods. Our proposed attack is a restricted attack method that aims to maliciously attack FR systems to expose more blind spots of them. It is not fair to compare our proposed method with unrestricted attacks that do not limit the magnitude of the adversarial perturbations. Therefore, we choose restricted attacks on FR that aim to maliciously attack FR systems [70] [72] [29] and state-of-the-art transfer attacks [60] [33] [54] [51] as our baseline.
Table 2: Comparisons of ASR (%) with multi-task attacks on LFW dataset. Models in the second row are victim models.
ASRd ASRi
Attack IR152 FaceNet MF
Lagrangian 3.9 26.5 100.0 26.0
Lagrangian + ours 7.3 36.4 100.0 26.6
DA 11.0 35.6 99.1 37.4
DA + ours 17.5 44.9 99.4 37.8
4.2 Comparison Study
We compare our proposed attack method with the state-of-the-art attacks on multiple FR models and datasets. Several adversarial examples are illustrated in Fig. 5. The attack performance results are shown in Table 1. Table 1 illustrates that the incorporation of our proposed attack method significantly enhances the dodging ASR of adversarial attacks. It is worth noting that the average black-box impersonation ASRs of the baseline attacks in Table 1 also increase after integrating our proposed attack method. This demonstrates the effectiveness of our proposed method in improving the dodging attack performance while simultaneously maintaining the impersonation attack performance. Furthermore, we conducted a comparison between our proposed Adv-Pruning and multi-task attacks using MF as the surrogate model on LFW based on DI. For our proposed method, we choose the corresponding multi-task attack as the attack for both the Priming and Restoration stages. The dodging ASR and average black-box impersonation ASR results are shown in Table 2. Table 2 underscores the effectiveness of our method in enhancing the dodging performance of multi-task attacks while maintaining the impersonation performance. To further validate our proposed attack method on additional FR models, we selected SIA [54] as Baseline and IR152 as the surrogate model. The experimental settings are consistent with those described in Table 1. The dodging ASR across multiple FR models is demonstrated in Fig. 3. As depicted in Fig. 3, the dodging ASR improves on multiple FR models after integrating our proposed method, further confirming the effectiveness of our attack.
Figure 5: The Illustration of adversarial examples crafted by various attacks. First column: some attacker images. Last column: the corresponding victim images. The second to fifth columns exhibit the corresponding adversarial face examples crafted by BPFA, BPFA + ours, BSR, and BSR + ours, respectively.
In practical application scenarios, victims can employ adversarial robust models to defend against adversarial attacks. Consequently, it becomes crucial to evaluate the performance of adversarial attacks on these robust models. In this study, we generate adversarial examples on the LFW dataset using MF as the surrogate model and assess the performance of various attacks on the adversarial robust models. The results are presented in Fig. 4. The letters following the en dash represent the surrogate models, with ’I’, ’F’, and ’M’ corresponding to IR152adv, FaceNetadv, and MFadv, respectively. Fig. 4 illustrates that the inclusion of our proposed method leads to improvements in both dodging and impersonation performance. These results serve as evidence of the effectiveness of our proposed method on adversarial robust models.
JPEG compression is a widely adopted method for image compression during transmission, concurrently acting as a defense mechanism against adversarial examples. To assess the effectiveness of our proposed attack under JPEG compression, we utilize DI as the baseline attack and MF as the surrogate model, evaluating the attack performance on ArcFace and CurricularFace models with experimental settings consistent with those described in Table 1. The results are illustrated in Fig. 6. These results demonstrate that across varying levels of JPEG compression, our proposed attack method consistently outperforms the baseline attack, thereby highlighting its effectiveness under JPEG compression.
The experimental results on negative cosine similarity loss, and Sibling-Attack are presented in the appendices.
4.3 Ablation Study
To delve into the properties of our proposed attack method, we conducted an ablation experiment using DI as the Baseline attack, with MF serving as the surrogate model on the LFW dataset. To confirm the effectiveness of our pruning method, we employed the Random Zeroing (RZ) method, which randomly sets adversarial perturbations to zero. We applied this method and our pruning method to free up 20% of the adversarial perturbations crafted by the Lagrangian attack. For Fine-tuning, we employed the Lagrangian attack as the method to further optimize the Lagrangian adversarial examples with a lower $\lambda$. The dodging attack ASR and average black-box impersonation ASR results are shown in Table 3. Table 3 demonstrates that our proposed pruning method for adversarial examples achieves a significantly smaller decrease in ASR than RZ after pruning 20% of adversarial perturbations, indicating the effectiveness of our pruning method. After being processed using the Pruning and Restoration stage of our proposed Adv-Pruning method, both impersonation and dodging ASR of the crafted adversarial face examples are recovered and higher than the Baseline attack method. These results demonstrate the effectiveness of our proposed Adv-Pruning in improving the dodging performance of adversarial attacks on FR without compromising the impersonation attack performance.
The sparsity ratio quantifies the proportion of adversarial perturbations that are allowed to be discarded during the Pruning stage. This ratio greatly impacts the performance of our proposed attack method. Hence, we conducted a sensitivity study on the sparsity ratio to analyze its effect on performance of the algorithm. The Lagrangian attack method based on DI is selected as the Baseline. We conduct a hyperparameter sensitivity study on LFW using FaceNet as the surrogate model, and adjust the value of $\tilde{\lambda}$ to ensure that the average black-box impersonation ASR results were within a 0.4% absolute difference compared to the Baseline. The dodging ASR results are illustrated in the right plot of Fig. 7. The results illustrate that the dodging ASR of our proposed method initially increases and then decreases as the sparsity ratio increases. When the sparsity ratio increases, a greater number of adversarial perturbations are pruned, creating more empty regions for the adversarial perturbations that favor dodging attacks in the Restoration stage. If the sparsity ratio is set to a too-high value, an excessive number of adversarial perturbations are allowed to be freed up, resulting in a degradation of performance for the adversarial examples crafted by the Priming stage. Consequently, the performance of adversarial face examples will decrease.
Figure 6: The dodging ASR under various JPEG Q values.
4.4 Analytical Study
Multi-identity Samples among the Natural Face Images: Multi-identity samples are intriguing samples that can be classified as multiple classes in FR. In this section, we will explore the existence of multi-identity samples among the natural face images.
Table 3: Comparisons of ASR (%) results of dodging attack and impersonation attacks on the LFW dataset. The models in the second row are the victim models.
ASRd ASRi
Attack IR152 FaceNet MF
Baseline 2.2 18.2 99.2 25.6
Lagrangian 3.9 26.5 100.0 26.0
Fine-tuning 4.2 26.3 100.0 25.6
RZ 0.6 4.4 94.8 15.1
Pruning 3.4 24.8 100.0 25.4
Adv-Pruning 5.4 32.5 100.0 26.3
We randomly select negative face pairs from the entire LFW dataset. Subsequently, we use MF as our FR model to extract the embeddings of the face images in each face pair and calculate the cosine similarity between the two images. If the cosine similarity surpasses the threshold, both images in the pair are classified as belonging to multiple identities, indicating that they are multi-identity samples. Our findings demonstrate the presence of multi-identity samples among the benign face images, as illustrated in the bottom left of Fig. 1. The multi-identity samples in Fig. 1 closely resemble the appearances of the identities they are classified into.
To analyze the cause of this phenomenon, we need to consider the properties of both the multi-identity samples and the FR model. Commonly-used FR models are well-trained and capable of correctly classifying the majority of benign face images. However, there are some benign face images that the FR model fails to classify accurately, and these samples are referred to as hard samples [21; 72]. Multi-identity samples are a specific type of hard sample known as hard negative samples. Typically, hard negative samples exhibit a similar appearance [67], indicating that the multi-identity samples among the benign face images share a resemblance.
Universality of Multi-identity Samples among Adversarial Face Examples: The previous impersonation methods craft adversarial face examples only use the impersonation loss $\mathcal{L}^{i}$. In this section, we will investigate the ratio of multi-identity samples and evaluate the effectiveness of previous impersonation methods in terms of dodging attacks. We utilize the Multi-identity Sample Ratio (MSR) to gauge the proportion of multi-identity samples in the adversarial face examples capable of executing successful impersonation attacks. These multi-identity samples can be recognized as both the attacker and victim identities in the setting of our paper. The detailed calculation method of MSR is in the appendices. We evaluate the MSR, impersonation ASR and dodging ASR using MF as the surrogate model on the LFW dataset and the results are demonstrated in the appendices. The results demonstrates that most of the crafted adversarial face examples are multi-identity samples in the black-box setting. This indicates that most adversarial face examples generated through previous impersonation attacks are unable to attain a successful dodging attack in the black-box setting. Nevertheless, in the white-box setting, the majority of adversarial face examples capable of executing successful impersonation attacks also demonstrate success in dodging attacks.
To analyze the reason, we consider the metric used to determine whether two face images belong to the same identity in FR. Since FR is an open-set task, we rely on the distance in the embedding space to make decisions. The top of Fig. 1 illustrates two decision boundaries for each FR model, one for attacker identity and one for victim identity. In the white-box setting, if we generate adversarial examples using $\mathcal{L}^{i}$, these adversarial examples can penetrate a space where they are recognized as the victim identity rather than the attacker identity. However, the decision boundary of the black-box model differs from that of the surrogate model. In the black-box setting, most adversarial examples are found between the decision boundaries of the black-box model, resulting in the majority of adversarial examples crafted using $\mathcal{L}^{i}$ being multi-identity samples. This demonstrates that the adversarial face examples are positioned near the decision boundary in the black-box setting.
The Trade-off Between the Impersonation Attacks and Dodging Attacks: Owing to the inherent conflict during the optimization process of impersonation and dodging losses in the black-box setting, there exists a trade-off between impersonation and dodging performance. We craft adversarial face examples using Lagrangian attack and our proposed Adv-Pruning on LFW based on DI. The average black-box results are demonstrated in the left plot of Fig. 7.
Figure 7: (a) The trade-off between the ASR (%) of impersonation attack and dodging attack of adversarial examples. (b) The dodging ASR (%) in different sparsity ratios.
The results illustrate that our proposed method can reduce the trade-off between impersonation and dodging performance in the black-box setting. The pruning operation of our proposed Adv-Pruning serves to sparsify the adversarial perturbations while preserving the impersonation performance. On the other hand, the restoration operation tends to introduce adversarial perturbations in the pruned areas, specifically favoring dodging attacks. These operations effectively enhance the dodging attack performance while maintaining the impersonation attack performance, ultimately mitigating the trade-off.
5 Conclusion
In this paper, we reveal the universality of multi-identity samples among adversarial face examples crafted by previous impersonation attacks and the success of an impersonation attack may not necessarily imply the success of dodging attacks on FR systems in the black-box setting. In order to improve dodging performance without compromising impersonation performance, we proposed a novel attack, namely Adv-Pruning. Adv-Pruning comprises Priming, Pruning, and Restoration Stages. Leveraging our proposed Adversarial Priority Quantification, we identify less prioritized adversarial perturbations with minimal impact on absolute model output variances. Through our proposed Biased Gradient Adaptation, biased gradient perturbations are applied to the sparsified regions, adapting adversarial face examples to a space favoring dodging attacks. Extensive experiments demonstrate the effectiveness of our proposed method.
Appendix A The Pseudo-code of Adversarial Pruning Attack Method
The pseudo-code of our proposed method based on the Lagrangian attack is illustrated in Algorithm 1.
Algorithm 1 Adversarial Pruning Attack Based on Lagrangian
@@ Appendix A The Pseudo-code of Adversarial Pruning Attack Method
0:  Negative face image pair $\{x^{s},x^{t}\}$, the number of the iterations of Priming stage $n$, the maximum number of iterations $m$, maximum allowable perturbation magnitude $\epsilon$, the surrogate FR model $\mathcal{F}$, the step size $\beta$ and $\gamma$.
@@
3:     $\mathcal{L}^{p}=\lambda\mathcal{L}^{i}+\mathcal{L}^{d}$triangleright Priming Stage
@@
Appendix B The Optimization Objective of the Impersonation and Dodging Attacks
The objective of impersonation attacks in our research is to craft an adversarial example $x^{adv}$ that causes the model $\mathcal{F}^{vct}$ to incorrectly classify it as the target sample $x^{t}$, while simultaneously maintaining a high degree of visual similarity between $x^{adv}$ and the original sample $x^{s}$. To be more specific, the objective can be expressed as follows:
where $\mathcal{D}$ refers to a predefined distance metric, while $\epsilon$ specifies the maximum magnitude of permissible perturbation.
In contrast, the dodging attacks introduced in this study aim to prevent the model $\mathcal{F}^{vct}$ from correctly recognizing $x^{adv}$ as the source sample $x^{s}$. This is achieved while maintaining a visually similar appearance between the adversarial example $x^{adv}$ and the original input $x^{s}$. In a manner similar to impersonation attacks, the objective of dodging attacks can be formulated as follows:
Appendix C The proof of the supremum of the absolute variance
In this section, we will provide the proof of the supremum of the absolute variance of the output of $\mathcal{F}$ before and after pruning. The absolute variance can be expressed as follows:
$\rho=\|\mathcal{F}\left(x+\delta\right)-\mathcal{F}\left(x\right)\|$ (15)